# Hermitian and non-Hermitian topology in active matter

### v1

@@ 1 Introduction
On another front, over the past few decades, the notion of topology has been widely applied to condensed matter physics [10, 11, 12]. Topology can tell us physical properties that are preserved under continuous deformations of parameters, such as the strength of interactions and geometry of the system. A representative example of topological properties is a genus of a manifold, which cannot be unchanged until we cut the manifold or stick its surfaces. Such a topological invariant distinguishes manifolds that can be continuously deformed into each other; for example, a donut can be deformed into a mug, while it cannot be deformed into a ball. One can consider defective structures like holes in spins and liquid crystals [13, 14], which are called topological defects and characterized by nontrivial topology around them. Since the topological properties are unchanged under the continuous deformations of parameters, such a topological defect is also robust against disorders.
In the field of condensed matter physics, topological defects in the wavenumber space, i.e., the band structures of electrons, are often discussed [15, 16]. Such nontrivial topology in band structures brings the notion of topological materials, which is exemplified by topological insulators. The first discovery of the topological band structure is the quantum Hall effect [17], where the Hall conductance is quantized. While the conventional quantum Hall effect is observed under the existence of the external magnetic field, the quantization of the Hall conductance only depends on the nontrivial topology of the band structure [18] and does not require the external magnetic field. In fact, by using the spin-orbit interaction, the quantum spin Hall effect, the time-reversal symmetric version of the quantum Hall effect, was proposed [19, 20], leading to the subsequent development of topological materials.
@@
The remaining part of this review is organized as follows. In Sec. 2, we introduce experimental setups of active matter. We also explain the hydrodynamic equation of active matter in Sec. 2.3, which is the starting point to discuss the active-matter counterparts of topological insulators. In Sec. 3, we introduce basic concepts of band topology in condensed matter physics. In particular, we discuss theoretical tools to analyze non-Hermitian topology, which is an emerging field dealing with the interplay between band topology and out-of-equilibrium physics. In Sec. 4, we review the studies on band topology in active matter. Finally, we discuss future perspectives in Sec. 5.
@@ 2.1.1 Flocking
Self-propelled particles are of prevalent interest in the field of statistical physics owing to their ability to collectively move in the same direction without a leader or direct aligning interactions. Animals, such as birds and fish, serve as typical examples of these particles [2], which can form flocks and schools at a collective scale. One early attempt to describe the collective dynamics of birds is the numerical simulation study of boids [37], which refers to bird-like (bird-oid) particles. Boids possess attractive, exclusive, and aligning interactions, allowing them to imitate the flocks of birds despite the symmetric role of the particles.
@@
$\theta_{i}(t+\Delta t)$ $=$ $\arg\left(\sum_{j\in U_{i}}e^{i\theta_{j}(t)}\right),$ (2)
where $\mathbf{r}_{i}(t)$ and $\theta_{i}(t)$ are the position and angle of the $i$th particle at time $t$. $\mathbf{e}_{\theta_{i}(t)}$ is a unit vector whose angle is $\theta_{i}(t)$. $U_{i}$ represents the vicinity of the $i$th particle, $U_{i}=\{\mathbf{r}^{\prime}|||\mathbf{r}^{\prime}-\mathbf{r}_{i}||<r_{0}\}$ with $r_{0}$ being a range of the aligning interaction. The Vicsek model exhibits collective motion akin to a flock of animals; even without attractive interactions as introduced in the boids, the active particles still formed dense clusters. Due to its simplicity, the Vicsek model and its variants [38, 39] have been widely used to analyze aspects of active systems as detailed later.
@@
Within the collective dynamics of living animals, a number of studies have been conducted on the natural swarm of midges [44, 45, 46]. The dynamical exponent, i.e., an exponent that relates the smarm size to the timescale of autocorrelation, has been explained by the dynamic renormalization procedure that incorporates inertia and friction in their latest model [47]. Other interesting examples of collective dynamics include the analysis of the dynamics of sheep [48], fish [49], and insects [50, 51], as well as human swarms [52], although their relevance to simple models of the Vicsek type is often unclear. The situation is presumably even more complex in social animals including ants and humans, which are also affected by social or psychological forces [53].
@@ 2.1.2 Cells and bacteria
The interaction between flagella and the substrate or boundaries can also affect the motion of bacteria. For example, the flagella can kick the wall during their rotation, which can change the direction of the bacteria. Thus, bacteria near the wall can show a curved motion, which leads to the chirality of active fluid [62].
The hydrodynamic interactions between bacteria can lead to the dynamical instability of the polar or nematic alignment, which can lead to chaotic bending flows observed in experiments of the collective motion of bacteria [3]. This bacterial turbulence, although an interesting phenomenon on its own, hinders the application of simple active fluid dynamics to predict the behavior of the collective dynamics of bacteria. Recent studies have proposed methods to control the active turbulence by using pillars or a liquid-crystal medium [63, 64]. Periodically located pillars can rectify the flow of active matter and the clockwise and anti-clockwise vortices can alternately appear in the squared regions surrounded by pillars. Another method uses rod-like passive particles, which will allow bacteria to align parallel or anti-parallel to the passive particles [65, 66].
@@
The nematic pattern can be easily characterized by the topological defects observed in the cell culture. Topological defects are defective structures around which the degree of the direction of active particles changes ±$\pm\pi$ or more, and typical defects called the +1/2 and -1/2 defects can only be observed when the cell interactions are nematic. The same topological defects have been observed in rod-like shaped cultured cells [67, 69, 70] as well as in epithelial cells [35], which are more isotropic and less bipolar at the single-cell level, yet can exhibit coordinated patterns that allow for defining the nematic order [71]. Defects in the crystal order of epithelial sheets in real animals have also been studied in Ref. [72].
@@ 2.1.3 Molecular motors
Active matter has been also studied at the molecular scale, which consists of components that are orders of magnitude smaller than cells and bacteria. Molecular motors [87] are typical examples of such microscopic active matter, which are abundant in cells. Specifically, two types of molecular motors and associated biological filaments are frequently used to study active matter dynamics; myosins and actin filaments [88], and kinesins and microtubules [89]. Myosins create a power of muscles by pulling filaments in opposite directions. On the other hand, kinesins walk on the microtubules and carry cargo that is attached to their heads. The kinesin’s walk can be regarded as an active particle on a one-dimensional chain and analyzed in some particle-based models, such as the one similar to the asymmetric exclusion process [32].
While molecular motors are prominent examples of active particles, filaments attached to these motors are also typically used to study the collective motion as active matter [36]. The motility assay is a typical setup of such experiments of active systems where molecular motors are fixed on a substrate and microtubules or actin filaments are free to move. This setup allows for the observation of the motor proteins’ ability to transport the filaments by moving their heads. From the coarse-grained view, the collective motion of such carried filaments is active rods. It is noteworthy that the motion of the filaments often becomes chiral, i.e., circular motions corresponding to the chirality of the molecules [36, 90].
@@ 2.1.4 Other systems
To study the dynamics of active matter, many studies have also utilized physical setups using non-biological active particles or agents. One straightforward way to realize active matter is using robots [95, 96]. Although the dynamics and interactions can be directly programmed in robots, there is often the problem of preparing them in macroscopic numbers since sophisticated robots tend to be expensive.
@@ 2.2 Biochemical network
where each component of $\mathbf{n}=(n_{1},n_{2},\cdots,n_{N})^{T}$ represents the concentration of a chemical species, and $\mathbf{f}(\mathbf{n})=(f_{1}(\mathbf{n}),f_{2}(\mathbf{n}),\cdots,f_{M}(\mathbf{n}))$ is the nonlinear vector function whose $j$th components are determined by the reaction rate of the $j$th chemical reaction. $S$ is a stoichiometric matrix whose $j$th row represents the numbers of molecules changed under the $j$th chemical reaction. In most cases, $f_{j}(\mathbf{n})$ is a nonlinear function of $\mathbf{n}$, while if the chemical reaction can be regarded as an effectively unimolecular reaction, $f_{j}(\mathbf{n})$ is proportional to the concentration of one chemical species $f_{j}(\mathbf{n})\propto n_{k}$. Therefore, assuming all the reactions can be regarded as unimolecular reactions, one can obtain a linear rate equation and construct an analogy between chemical reactions and quantum mechanics [104] (Fig. 2). In fact, a previous study [107] has discussed the topological origin of localization in chemical reactions as we will review below.
@@ 2.3.1 Toner-Tu model
Toner and Tu [112] also performed the renormalization analysis of the hydrodynamic equations. They first assumed a homogeneous ordered state $\rho=\rho_{0}$, $\mathbf{v}=v_{0}\mathbf{e}_{x}$, where $\mathbf{e}_{x}$ is the unit vector in the $x$ direction. Then, they considered the fluctuation from the ordered state, $\delta\rho=\rho-\rho_{0}$ and $\delta\mathbf{v}=(\sqrt{1-v_{\perp}^{2}},v_{\perp})$, where $v_{\perp}$ is the velocity perpendicular to the direction of the steady-state velocity. The hydrodynamic equations can be modified into the equations of $\rho$ and the scalar version of $v_{\perp}$ (denoted as $h$) as follows:
@@
Numerical studies of the hydrodynamic equations have also reproduced the nonequilibrium features of active matter [114, 115]. Specifically, the active hydrodynamics exhibits a flocking-like behavior, i.e., ordered cluster formation. Furthermore, phase separations and band formulations observed in Vicsek-type models are also observed in hydrodynamics.
Nevertheless, we mention that the hydrodynamic equations are not perfect; they fail to describe some features observed in experiments or even in numerical simulations of particle models. For example, the Vicsek model can selectively show micro-phase separation with band-shaped aggregations with a constant width [39, 116]. However, the hydrodynamics cannot explain the selectivity of the width. In more realistic situations, active fluids often show chaotic dynamics called active turbulence [34]. To analyze such turbulence-like phenomena, previous research has phenomenologically introduced a higher-order term of the Swift-Hohenberg-type [3, 34, 117] that is proportional to the fourth-order derivative of the velocity field, $(\nabla^{2})^{2}\mathbf{v}$.
@@ 2.3.2 From micro to hydrodynamic models
[…]
×$\times f(\mathbf{r},\theta_{1},t)f(\mathbf{r},\theta_{2},t)\delta_{m\pi}(\Phi(\theta_{1},\theta_{2})+\eta-\theta)$
@@
[…]
@@
are equivalent to the hydrodynamic equations of active matter.
@@ 2.3.3 Hydrodynamics of various self-propelled particles
While the pioneering work by Vicsek and coworkers studied self-propelled particles with polar (i.e., ferromagnetic) interactions, active matter can have various symmetries in its interaction and motion. A commonly appearing example is the self-propelled rods [110], which move unidirectionally but with nematic interactions, favoring both parallel and antiparallel states. These self-propelled rods can model the motion of, e.g., bacteria. Another type of active matter is active nematics [85], which are typically seen in cultured cells that undergo flipping in their direction of motion.
@@
where $\theta$ is the mean orientation of the particles and $S$ represents the local nematic order. This tensor has two degrees of freedom (diagonal and off-diagonal terms), and these two components and the density are the slow variables in the hydrodynamics of active nematics. We note that these nematic order terms correspond to the second Fourier components in the Boltzmann equation (8). In active nematics, the symmetry prohibits the relevant velocity field and thus the hydrodynamics includes only the density and the nematic tensor. In self-propelled rods, both the velocity field and the nematic tensor are leading-order terms, and the hydrodynamic equation becomes
@@
$+\tau||\mathbf{P}||^{2}Q-\chi\nabla\cdot(\mathbf{P}\tilde{Q})-\kappa\mathbf{P}\cdot(\nabla\tilde{Q}),$ (21)
@@
Classification of active matter can be done from another point of view, namely, “dry” or “wet” active matter [7, 110]. A dry system refers to active particles that involve none or negligible hydrodynamic interactions (even if they are surrounded by fluids), which is mostly the case for cultured cells. A wet system, on the other hand, refers to a case where hydrodynamic interactions are significant, such as in bacterial turbulence. So far, we have mainly discussed the hydrodynamics of dry active matter, while the hydrodynamics of wet active matter has been also derived in Ref. [7]. Such hydrodynamic equations include the active stress tensor that the self-propulsion of each active particle generates, which can be approximated by a force dipole (Stokeslet) far from each particle [125, 126]. Adding such hydrodynamic interaction makes the dynamics of wet active matter more difficult to analyze and calculate, but can reproduce more realistic behaviors in wet active matter [127].
Active matter can also exhibit rotating motion, which further breaks the left-right symmetry of particles. Such chiral active matter can be found in various situations, such as bacteria [62] and molecular motors [90]. In hydrodynamics, the chirality of active particles adds two relevant terms, a Coriolis-like force and odd viscosity [8, 128]. The Coriolis-like force is described as $\omega_{0}\nabla\times\mathbf{v}$, where $\omega_{0}$ is determined from the frequency of the self-rotation of the chiral active particles. This force acts on active matter perpendicularly to the velocity field in a similar manner to the conventional Coriolis force. The other chirality term, odd viscosity $\nu^{o}\nabla^{2}\nabla\times\mathbf{v}$, is an unconventional viscosity that is antisymmetric under the time reversal. This term does not lead to energy dissipation but induces the Hall-like linear response perpendicular to the pressure. Both Coriolis-like force and odd viscosity play an essential role in topological active matter as discussed in Sec. 4.2.
3 Hermitian and non-Hermitian band topology

### v2

@@ 1 Introduction
On another front, over the past few decades, the notion of topology has been widely applied to condensed matter physics [10, 11, 12]. Topology can tell us physical properties that are preserved under continuous deformations of parameters, such as the strength of interactions and geometry of the system. A representative example of topological properties is a genus of a manifold, which cannot be unchanged until we cut the manifold or stick its surfaces. Such a topological invariant distinguishes manifolds that can be continuously deformed into each other; for example, a donut can be deformed into a mug, while it cannot be deformed into a ball. Similar topology has been found in real-space configurations of spins and liquid crystals [13, 14], where topological charges are defined around their defective structures.
In the field of condensed matter physics, a different kind of topology has also been discussed in the wavenumber space, i.e., the band structures of electrons. Such nontrivial topology in band structures brings the notion of topological materials, which is exemplified by topological insulators. The first discovery of the topological band structure is the quantum Hall effect [15], where the Hall conductance is quantized. While the conventional quantum Hall effect is observed under the existence of the external magnetic field, the quantized Hall conductance is proportional to the topological number (Chern number) of occupied bands [16], which can be nonzero even without external magnetic fields by instead using internal magnetic structures. Furthermore, symmetries enrich the topological phases of matter. In fact, by using the spin-orbit interaction, the quantum spin Hall effect, the time-reversal symmetric version of the quantum Hall effect, was proposed [17, 18], which stimulates the subsequent studies on topological materials.
@@
The remaining part of this review is organized as follows. In Sec. 2, we introduce experimental setups of active matter. We also explain the hydrodynamic equation of active matter in Sec. 2.3, which is the starting point to discuss the active-matter counterparts of topological insulators. In Sec. 3, we introduce basic concepts of band topology in condensed matter physics. In Sec. 4, we introduce topological phenomena unique to non-Hermitian systems, which is an emerging field dealing with the interplay between band topology and out-of-equilibrium physics. We also provide theoretical tools to analyze non-Hermitian topology called the non-Bloch band theory. In Sec. 5, we review the studies on band topology in active matter. We also explain how to combine classical physics and topological physics explained in the above sections via the linearization of active hydrodynamics. Finally, we discuss future perspectives in Sec. 6.
@@ 2.1.1 Flocking
Self-propelled particles are of prevalent interest in the field of statistical physics owing to their ability to collectively move in the same direction without a leader or direct aligning interactions. Animals, such as birds and fish, serve as typical examples of these particles [2], which can form flocks and schools at a collective scale (cf. Fig. 1(a)). One early attempt to describe the collective dynamics of birds is the numerical simulation study of boids [35], which refers to bird-like (bird-oid) particles. Boids possess attractive, exclusive, and aligning interactions, allowing them to imitate the flocks of birds despite the symmetric role of the particles.
@@
$\theta_{i}(t+\Delta t)$ $=$ $\arg\left(\sum_{j\in U_{i}}e^{i\theta_{j}(t)}\right)+\xi_{i}(t),$ (2)
[…]
@@
Within the collective dynamics of living animals, a number of studies have been conducted on the natural swarm of midges [44, 45, 46]. The dynamical exponent, i.e., an exponent that relates the smarm size to the timescale of autocorrelation, has been explained by the dynamic renormalization procedure that incorporates inertia and friction in the latest model by Cavagna et al. [47]. Other interesting examples of collective dynamics include the analysis of the dynamics of sheep [48], fish [49], and insects [50, 51], as well as human swarms [52], although their relevance to simple models of the Vicsek type is often unclear. The situation is presumably even more complex in social animals including ants and humans, which are also affected by social or psychological forces [53].
@@ 2.1.2 Cells and bacteria
The interaction between flagella and the substrate or boundaries can also affect the motion of bacteria. In particular, collision and hydrodynamic interaction between rotating flagella and wall can induce chirality of active fluid, i.e., bacteria near the wall show a curved motion [62, 63]. Hydrodynamic interaction between pusher bacteria like E. coli and a wall stabilizes swimming in a parallel direction to the wall [64]. Then, the rotating body and flagella further induce the hydrodynamic interaction that leads to the chiral bacterial motion. At the same time, fluctuating dynamics of bacteria induce collisions with the wall, which is also a source of chiral motions of the bacteria. Experimental studies (such as Ref. [65]) imply that the collision may be a major contribution to the chiral motion rather than the hydrodynamic interaction.
The hydrodynamic interactions between bacteria can lead to the dynamical instability of the polar or nematic alignment, which can lead to chaotic bending flows observed in experiments of the collective motion of bacteria (cf. Fig. 1(b)) [3]. This bacterial turbulence, although an interesting phenomenon on its own, hinders the application of simple active fluid dynamics to predict the behavior of the collective dynamics of bacteria. Recent studies have proposed methods to control the active turbulence by using pillars or a liquid-crystal medium [66, 67]. Periodically located pillars can rectify the flow of active matter and the clockwise and anti-clockwise vortices can alternately appear in the squared regions surrounded by pillars. If one uses other periodic alignments of pillars that have sublattice structures, the rectified flow can break symmetries and enable the realization of an active-matter counterpart of topological insulators discussed in Sec. 5.2.2. Another method uses rod-like passive particles, which will allow bacteria to align parallel or anti-parallel to the passive particles [68, 69].
@@
The nematic pattern can be easily characterized by the topological defects observed in the cell culture. Topological defects are defective structures around which the degree of the direction of active particles changes $\pm\pi$ or more, and thus at the center, one cannot determine the direction. Since the rotation of the direction of active particles must be quantized by $\pi$, we cannot continuously remove this defect (i.e., no rotations of directions of nematics), which leads to the long-term stability of the defect with a topological origin. Typical defects called the +1/2 and -1/2 defects, where around $n/2$ defects, the direction of active particles changes $n\pi$, can only be observed when the cell interactions are nematic. The same topological defects have been observed in rod-like shaped cultured cells [70, 72, 73] as well as in epithelial cells [33] (cf. Fig. 1(c)), which are more isotropic and less bipolar at the single-cell level, yet can exhibit coordinated patterns that allow for defining the nematic order [74]. Defects in the crystal order of epithelial sheets in real animals have also been studied in Ref. [75].
@@ 2.1.3 Molecular motors
Active matter has been also studied at the molecular scale, which consists of components that are orders of magnitude smaller than cells and bacteria. Molecular motors [90] are typical examples of such microscopic active matter, which are abundant in cells. Specifically, two types of molecular motors and associated biological filaments are frequently used to study active matter dynamics; myosins and actin filaments [91], and kinesins and microtubules [92]. Myosins create a power of muscles by pulling filaments in opposite directions. On the other hand, kinesins walk on the microtubules and carry cargo that is attached to their heads. The kinesin’s walk can be regarded as an active particle on a one-dimensional chain and analyzed in some particle-based models, such as the one similar to the asymmetric exclusion process [30], which can be utilized as a typical setup to study non-Hermitian topology discussed in Sec. 3.
While molecular motors are prominent examples of active particles, filaments attached to these motors are also typically used to study the collective motion as active matter [34]. The motility assay is a typical setup of such experiments of active systems where molecular motors are fixed on a substrate and microtubules or actin filaments are free to move (cf. Fig. 1(d)). This setup allows for the observation of the motor proteins’ ability to transport the filaments by moving their heads. From the coarse-grained view, the collective motion of such carried filaments is active rods. It is noteworthy that the motion of the filaments often becomes chiral, i.e., circular motions corresponding to the chirality of the molecules [34, 93].
@@ 2.1.4 Other systems
To study the dynamics of active matter, many studies have also utilized physical setups using non-biological active particles or agents. One straightforward way to realize active matter is using robots [98, 99]. Although the dynamics and interactions can be directly programmed in robots, there is often the problem of preparing them in macroscopic numbers since sophisticated robots tend to be expensive. However, they are still useful to examine topological physics as discussed in Sec. 5.3.1, because topological phenomena tend to be robust against the finite-size effect and seen in a small system.
@@ 2.2 Biochemical network
where each component of $\mathbf{n}=(n_{1},n_{2},\cdots,n_{N})^{T}$ represents the concentration of a chemical species, and $\mathbf{f}(\mathbf{n})=(f_{1}(\mathbf{n}),f_{2}(\mathbf{n}),\cdots,f_{M}(\mathbf{n}))$ is the nonlinear vector function whose $j$th components are determined by the reaction rate of the $j$th chemical reaction. $S$ is a stoichiometric matrix whose $j$th row represents the numbers of molecules changed under the $j$th chemical reaction. In most cases, $f_{j}(\mathbf{n})$ is a nonlinear function of $\mathbf{n}$, while if the chemical reaction can be regarded as an effectively unimolecular reaction, $f_{j}(\mathbf{n})$ is proportional to the concentration of one chemical species $f_{j}(\mathbf{n})\propto n_{k}$. Therefore, assuming all the reactions can be regarded as unimolecular reactions, one can obtain a linear rate equation and construct an analogy between chemical reactions and quantum mechanics [107] (Fig. 2). In fact, a previous study [110] has discussed the topological origin of localization in chemical reactions as we will review in Sec. 5.4.
@@ 2.3.1 Toner-Tu model
Toner and Tu [115] also performed the renormalization analysis of the hydrodynamic equations. They first assumed a homogeneous ordered state $\rho=\rho_{0}$, $\mathbf{v}=v_{0}\mathbf{e}_{x}$, where $\mathbf{e}_{x}$ is the unit vector in the $x$ direction. Then, they considered the fluctuation from the ordered state, $\delta\rho=\rho-\rho_{0}$ and $\delta\mathbf{v}=(\sqrt{v_{0}^{2}-v_{\perp}^{2}}-v_{0},v_{\perp})$, where $v_{\perp}$ is the velocity perpendicular to the direction of the steady-state velocity. The hydrodynamic equations can be modified into the equations of $\rho$ and the scalar version of $v_{\perp}/v_{0}$ (denoted as $h$) as follows:
@@
Numerical studies of the hydrodynamic equations have also reproduced the nonequilibrium features of active matter [117, 118]. Specifically, the active hydrodynamics exhibits a flocking-like behavior, i.e., ordered cluster formation. Furthermore, phase separations and band formulations observed in Vicsek-type models are also observed in hydrodynamics. Based on these theoretical and numerical results, the Toner-Tu equations are believed to capture some universal properties of active matter and widely used to study the phases and dynamics of active matter. In particular, most of the theoretical proposals of topological active matter are based on the Toner-Tu equations or their variants, as is discussed in Sec. 5.
Nevertheless, we mention that the hydrodynamic equations are not perfect; they fail to describe some features observed in experiments or even in numerical simulations of particle models. For example, the Vicsek model can selectively show micro-phase separation with band-shaped aggregations with a constant width [39, 119]. However, the hydrodynamics cannot explain the selectivity of the width. In more realistic situations, active fluids often show chaotic dynamics called active turbulence [32]. To analyze such turbulence-like phenomena, previous research has phenomenologically introduced a higher-order term of the Swift-Hohenberg-type [3, 32, 120] that is proportional to the fourth-order derivative of the velocity field, $(\nabla^{2})^{2}\mathbf{v}$. The effect of this correction term has not been investigated in topological active matter and thus should be an important issue to be solved for discussing topology of real biological active matter.
@@ 2.3.2 From micro to hydrodynamic models
Another advantage of this derivation is that one can extend this method to any types of active-particle models and thus can systematically derive the corresponding hydrodynamics even if the model is too complicated to phenomenologically specify the leading-order terms. In fact, a previous study on topological active matter [124] derived the hydrodynamics from a particle model and then discussed its connection to topological physics, which we introduce later in Sec. 5.3.2.
@@
$-\lambda f+\lambda\int d\theta^{\prime}\int d\eta P_{\sigma}(\eta)\delta_{m\pi}(\theta^{\prime}+\eta-\theta)f(\mathbf{r},\theta^{\prime},t),$
$I_{col}[f]$ $=$ $-f(\mathbf{r},\theta,t)\int d\theta^{\prime}K(\theta^{\prime}-\theta)f(\mathbf{r},\theta^{\prime},t)$
$+\int d\theta_{1}\int d\theta_{2}\int d\eta P_{\sigma}(\eta)K(\theta_{2}-\theta_{1})$
$\times f(\mathbf{r},\theta_{1},t)f(\mathbf{r},\theta_{2},t)\delta_{m\pi}(\Phi(\theta_{1},\theta_{2})+\eta-\theta)$
@@
$I_{k,q}=(1/2\pi)\int d\Delta K(\Delta)e^{-iq\Delta+ikH(\Delta)},$ (13)
and $P_{k}=\int d\eta P_{\sigma}(\eta)e^{ik\eta}$. We also use $\nabla$ and $\nabla^{\ast}$ to represent the spatial derivative, $\nabla f=\partial_{x}f+i\partial_{y}f$ and $\nabla^{\ast}f=\partial_{x}f-i\partial_{y}f$. Finally, keeping the relevant terms, the equations corresponding to $k=0$ and $k=1$,
@@
are equivalent to the hydrodynamic equations of dry aligning dilute active matter (see the next section for the definition of dry active matter).
@@ 2.3.3 Hydrodynamics of various self-propelled particles
While the pioneering work by Vicsek and coworkers studied self-propelled particles with polar (i.e., ferromagnetic) interactions, active matter can have various symmetries in its interaction and motion as shown in Fig. 3(a). A commonly appearing example is the self-propelled rods [113], which move unidirectionally but with nematic interactions, favoring both parallel and antiparallel states. These self-propelled rods can model the motion of, e.g., bacteria. Another type of active matter is active nematics [88], which are typically seen in cultured cells that undergo flipping in their direction of motion.
@@
where $\theta$ is the mean orientation of the particles and $S$ represents the local nematic order. This tensor has two degrees of freedom, $S$ and $\theta$, and these two components and the density are the slow variables in the hydrodynamics of active nematics. We note that these nematic order terms correspond to the second Fourier components in the Boltzmann equation (8). In active nematics, the symmetry prohibits the relevant velocity field and thus the hydrodynamics includes only the density and the nematic tensor. In self-propelled rods, both the velocity field and the nematic tensor are leading-order terms, and the hydrodynamic equation becomes
@@
[…]
Classification of active matter can be done from another point of view, namely, “dry” or “wet” active matter [7, 113]. A dry system refers to active particles that break momentum conservation by, e.g., ignoring hydrodynamic interactions via fluids surrounding active particles, which is mostly the case for cultured cells. A wet system, on the other hand, refers to a case where hydrodynamic interactions are significant and thus one needs to consider momentum-conserving hydrodynamics, such as in bacterial turbulence. So far, we have mainly discussed the hydrodynamics of dry active matter, while the hydrodynamics of wet active matter has been also derived in Ref. [7]. Such hydrodynamic equations include the active stress tensor that the self-propulsion of each active particle generates, which can be approximated by a force dipole (Stokeslet) far from each particle [65, 129]. Adding such hydrodynamic interaction makes the dynamics of wet active matter more difficult to analyze and calculate, but can reproduce more realistic behaviors in wet active matter [130].
Active matter can also exhibit rotating motion, which further breaks the left-right symmetry of particles (cf. Fig. 3(b)). Such chiral active matter can be found in various situations, such as bacteria [62] and molecular motors [93] (cf. Secs. 2.1.2 and 2.1.3). In hydrodynamics, the chirality of active particles adds two relevant terms, a Coriolis-like force and odd viscosity [8, 131]. The Coriolis-like force is described as $\omega_{0}\nabla\times\mathbf{v}$, where $\omega_{0}$ is determined from the frequency of the self-rotation of the chiral active particles. This force acts on active matter perpendicularly to the velocity field in a similar manner to the conventional Coriolis force. The other chirality term, odd viscosity $\nu^{o}\nabla^{2}\nabla\times\mathbf{v}$, is an unconventional viscosity that is antisymmetric under the time reversal. This term does not lead to energy dissipation but induces the Hall-like linear response perpendicular to the pressure. Both Coriolis-like force and odd viscosity play an essential role in topological active matter as discussed in Sec. 5.2.
3 Hermitian band topology